OpenStudy (anonymous):

What is wrong with solving the equation x2=3x by dividing both sides of the equation by x? Explain the correct method for solving this equation.

OpenStudy (anonymous):

because you are not allowed to divide by 0 if \(x=0\) you have a solution because you would get \(0=0\)

OpenStudy (anonymous):

correct method is to think "if twice a number is the same as three times a number, that number better be 0"

OpenStudy (anonymous):

well actually correct method would be two write \[x^2-2x=0\] \[x(x-2)=0\] etc

OpenStudy (anonymous):

THANKS SATELLITE GREAT REACTION TIME

OpenStudy (anonymous):

ignore my second answer i read it incorrectly i thought it said \(2x=3x\) but actually it was \(x^2=3x\) so all is wrong

OpenStudy (anonymous):

let me start again and do it right

OpenStudy (anonymous):

\[x^2=3x\] \[x^2-3x=0\] \[x(x-3)=0\] \[x=0\] or \[x=3\]